OpenStudy (anonymous):

Solve for x: 2/5 (x − 2) = 4x.

OpenStudy (anonymous):

help me plz

OpenStudy (texaschic101):

2/5 (x - 2) = 4x --- distribute through the parenthesis 2/5x - 4/5 = 4x --- if you do not want to mess with fractions, multiply everything by common denominator of 5 2x - 4 = 20x --- subtract 2x from both sides -4 = 20x - 2x -4 = 18x -- divide by 18 on both sides -4/18 = x -- reduce - 2/9 = x or with fractions.. 2/5 (x - 2) = 4x 2/5x - 4/5 = 4x -- subtract 2/5x -4/5 = 4x - 2/5x -4/5 = 20/5x - 2/5x -4/5 = 18/5x -- now divide by 18/5 (-4/5) / (18/5) = x -4/5 * 5/18 = x - 20/90 = x -- reduce -2/9 = x

OpenStudy (texaschic101):

questions ?

OpenStudy (anonymous):

wow thanks you

OpenStudy (texaschic101):

anytime :)